# Reflection on Code Contributor Demographics and Collaboration Patterns in the Rust Community

### v2

@@ Abstract.
Open-source software communities thrive on global collaboration and contributions from diverse participants. This study explores the Rust programming language ecosystem to understand its contributors’ demographic composition and interaction patterns. Our objective is to investigate the phenomenon of participation inequality in key Rust projects and the presence of diversity among them. We studied GitHub pull request data from the year leading up to the release of the latest completed Rust community annual survey in 2023. Specifically, we extracted information from three leading repositories: Rust, Rust Analyzer, and Cargo, and used social network graphs to visualize the interactions and identify central contributors and sub-communities. Social network analysis has shown concerning disparities in gender and geographic representation among contributors who play pivotal roles in collaboration networks and the presence of varying diversity levels in the sub-communities formed. These results suggest that while the Rust community is globally active, the contributor base does not fully reflect the diversity of the wider user community. We conclude that there is a need for more inclusive practices to encourage broader participation and ensure that the contributor base aligns more closely with the diverse global community that utilizes Rust.
Social networks, OSS communities, Rust, DEI
@@ 1. Introduction
Open-source software (OSS) communities form the backbone of contemporary software development, enabling global collaboration and innovation. Participants from diverse backgrounds contribute in various ways (i.e., reporting issues, submitting pull requests (PRs), conducting code reviews, and updating documentation). These participants often contribute in their free time and at a level that aligns with their expertise and availability, fostering an environment of collective growth and learning (Lakhani and Wolf, 2005; Gerosa et al., 2021). Despite the seemingly vast pool of available developers, research shows that much of OSS development is driven by a small percentage of individuals. This phenomenon, known as “participation inequality” (Gasparini et al., 2020a), highlights the disparity in contribution levels, with most work concentrated among a few core contributors. Mature OSS communities, such as the Apache Software Foundation  [^1] https://www.apache.org/ are known to be inclusive (Feng et al., 2023) but have been plagued by the same issue (Chelkowski et al., 2016).
@@
Like many other OSS projects, the Rust community is organized into a structured membership system (Toral et al., 2010). Contributors—those who consistently contribute to the core Rust projects—play a pivotal role in its development. Full members of the Rust GitHub organization, who take on leadership roles such as voting on Requests for Comments (RFCs), help shape the project’s strategic direction.  [^3] https://forge.rust-lang.org/compiler/membership.html However, the extent to which these contributors reflect the global diversity of Rust’s user base remains underexplored. This disconnect could greatly impact the community’s ability to foster an inclusive environment that attracts diverse contributors worldwide.
@@
  • RQ3: To what extent do sub-communities and top contributors within the Rust ecosystem exhibit diversity?
To answer these research questions, we focused on the core Rust compiler, Rust Analyzer (an integrated development environment (IDE) front-end), and Cargo (the Rust package manager). We examined the contributions and interactions within these repositories. Utilizing the GitHub API, we extracted PRs, authors, and reviewers. Then, we enriched this dataset with demographic information such as geographic location and likely gender. We then constructed social network graphs in R, applying the Louvain community detection algorithm to identify sub-communities within the network. Diversity within these sub-communities was quantitatively assessed using Blau’s Index. At the same time, various network metrics were calculated to evaluate the structure and dynamics of Rust contributor interactions.
Our findings revealed that the diversity of active contributors does not fully reflect the broader Rust user community’s demographic composition. The contributor base is heavily concentrated in specific geographic regions, with a predominance of contributors from the United States and Germany. Gender diversity remains low, with men contributors dominating participation. Additionally, non-member contributors play a crucial role in the collaboration networks, often forming distinct clusters independent of official Rust organization members. Sub-communities demonstrate substantial geographic diversity; however, gender diversity remains notably low across all groups. These findings suggest that the community may need more inclusive practices to attract a broader range of participants to reflect its user base as it leads to more adoption, which in turn leads to more job creation that benefits the whole ecosystem (Carter and Groopman, 2021), aligns with Rust’s ethos of empowering everyone and capturing contributors from upcoming communities in Southern and Eastern parts of the world (GitHub, 2023).
@@ 2. Related work
The success and sustainability of OSS projects are intrinsically linked to their contributor communities’ collaborative efforts and network structures. Previous research has focused on understanding these dynamics, particularly through the lens of social network analysis and diversity studies.
@@ 2.1. Social Network Analysis in OSS Communities
An empirical analysis of GitHub project social team structures reveals a significant correlation between the speed of pull-request processing and the density of the collaboration network (Mezouar et al., 2019). The study found that smaller to medium-sized projects benefit from a tightly-knit core of maintainers who engage in repeated interactions, facilitating more efficient review and integration of contributions. Building on these findings, we aim to investigate similar dynamics within three significant repositories in the Rust ecosystem. Our analysis will examine the network structures between members (typically those with integration and approval access) and non-members, further exploring how geographic and gender demographics influence interaction patterns.
@@ 2.3. Rust Community
Recent works have used network analysis (Imtiaz et al., 2023) to the Rust ecosystem, constructing developer networks based on collaboration activities. The study proposed integrating contributor reputation metrics to enhance dependency management practices, highlighting the importance of social factors in technical decision-making processes within OSS communities.
Our study seeks to bridge the current state of research done in the Rust ecosystem gap by explicitly analyzing the social dynamics and demographic diversity within the Rust community. By integrating social network analysis with demographic data such as geographic location and gender, we aim to uncover how diversity manifests within sub-communities and influences collaboration patterns. Unlike previous studies that have focused solely on centrality measures or diversity aspects, our research provides a comprehensive view combining both elements, offering deeper insights into the factors that drive participation and collaboration in the Rust ecosystem.
@@ 3.1. Projects Selection Criteria
As of January 2025, the Rust organization on GitHub comprises approximately 222 repositories. To focus our analysis on the most impactful projects, we selected the top three repositories based on the number of stars—a common indicator of popularity and community engagement (Ray et al., 2017). We excluded repositories, such as book, mdBook, and rustlings, as they are intended for documentation and learning purposes. After this refinement, our study concentrated on three core repositories:
@@
Furthermore, these three repositories represent distinct facets of the Rust ecosystem, each demanding unique skill sets and domain expertise. The Rust compiler involves low-level systems programming and compiler design, requiring contributors to understand language internals and performance optimization deeply. Rust Analyzer focuses on IDE integration and developer tooling, necessitating knowledge in language server protocols, code analysis, and user experience considerations. Cargo, as the package manager, revolves around dependency management and building systems, where contributors need proficiency in software distribution and project configuration. Each project follows unique development and collaboration practices tailored to the specific goals and challenges mentioned in their dedicated guidelines. [^5] https://rustc-dev-guide.rust-lang.org/contributing.html [^6] https://github.com/rust-lang/rust-analyzer/blob/master/docs/dev/guide.md [^7] https://doc.crates.io/contrib/ By including these repositories, our study captures a comprehensive range of contribution dynamics and interactions within the Rust community, providing insights into how varying technical domains influence participation and collaboration dynamics.
@@ 3.2. Extracting PRs, Author, and Reviewers
We posit that individuals who invested the time to complete the comprehensive Rust Annual Community Survey 2023 (Team, 2024)—which takes approximately 25 minutes—represent the community’s distribution due to their significant level of engagement and commitment. Hence, to understand and compare the contributors’ dynamics preceding the Rust Annual Survey 2023, we collected PR data from the year prior to the survey’s release, specifically from December 18th, 2022, to December 18th, 2023.
@@
Additionally, we excluded PRs involving automated bots such as dependabot, bors, and rustbot to ensure our analysis focused solely on human contributors.
3.3. Enriching data with Full Name, Location and Membership
@@
To ascertain the geographical distribution of contributors, we analyzed the location details provided in user profiles. We utilized the geograpy3888https://pypi.org/project/geograpy3/ Python package, which leverages the Natural Language Toolkit (NLTK) framework, to parse the city or location strings and extract corresponding country information. The extracted data were then converted to ISO2 country codes [^9] https://www.iso.org/obp/ui/#search using the country_converter package. To ensure the highest accuracy, we manually verified the location data by searching the location details (city, country emoji, etc) online with the assigned country code, correcting any inconsistencies or errors identified during the automated extraction process. Overall, about a third of the data was inconsistently classified. For example, ”Earth” and ”UK” were being classified under the ”US” country code.
Here’s the table with all values rounded to whole numbers (no decimal points):
Table 1. Statistics of code contributors demographics in three crucial Rust projects
@@
For gender identification, we relied on contributors’ full names. We employed Stanford’s Named Entity Recognition (NER) tool, available through the Python package stanza [^10] https://stanfordnlp.github.io/stanza/ to determine whether the extracted names referred to persons. This filtered out usernames that might represent organizations, projects, or non-person entities. Only names recognized as ”PERSON” entities were considered in subsequent analyses. Subsequently, we used the NamSor API, a widely recognized name classification tool in academic research (Sebo, 2021; Qiu et al., 2019), to predict the likely gender of each contributor. By inputting the full name and corresponding ISO2 country code into NamSor, we obtained gender probabilities for each individual. To maintain a high level of confidence in our gender assignments, we included only those contributors for whom NamSor provided a gender probability of 80% or higher. Any cases with lower confidence levels or potential false positives underwent additional manual verification. Both the authors independently reviewed these ambiguous cases, and only upon mutual agreement were the final gender classifications assigned.
@@ 3.4. Social Network Analysis
Post enriched data collection, we constructed social network graphs using the R programming language and the igraph [^11] https://igraph.org/ package. Each contributor is represented as a node, and edges represent user interactions in these graphs. Specifically, we constructed two types of graphs, directed graphs where an edge from the author node to the reviewer node signifies that the author submitted a PR that was reviewed by the reviewer, indicating the directionality of the collaboration. Additionally, undirected graphs were considered where edges indicate that two users have interacted through PRs without implying a specific direction. To identify clusters within the social network, we applied the Louvain community detection algorithm, which efficiently uncovers communities by optimizing modularity (Blondel et al., 2008).
This algorithm allowed us to detect sub-communities where contributors are more densely connected among themselves than with the rest of the network. Incorporating demographic details into our analysis enabled us to understand various collaboration dynamics. We examined how members and non-members of the Rust GitHub organization collaborate, exploring whether official affiliation influences interaction patterns. Additionally, we assessed the diversity within sub-communities by considering contributors’ geographic locations (using ISO2 country codes) and likely gender. This approach provided insights into how contributors from different backgrounds interact within the network.
The graphs were constructed by including nodes only when the respective node demographic information was available. To ensure the accuracy and relevance of our network analysis, we removed self-loops—edges that connect nodes to themselves—as they do not represent meaningful collaborations between different contributors. Furthermore, we excluded isolated nodes that remained disconnected after the removal of self-loops, focusing our analysis on the interconnected portion of the network where actual interactions occur.
@@
To capture sub-community geographic diversity, we constructed graphs using ISO2 country codes and used Blau’s Index(Blau, 1977) to calculate the diversity quantitatively in identified sub-communities across the projects (Shameer et al., 2023).
@@
A higher Blau’s Index value indicates greater diversity within the sub-community, while a lower value suggests homogeneity. By applying this metric to the geographic distributions of contributors, we were able to quantitatively compare the diversity levels across different sub-communities within the Rust ecosystem. Additionally, we employed centrality measures (Newman, 2018)—such as degree centrality (the number of direct connections a node has), betweenness centrality (the extent to which a node lies on paths between other nodes), and eigenvector centrality (the influence of a node based on the importance of its connections)——to identify influential contributors within the network. By examining the demographics of these central contributors, we aimed to understand which user groups hold significant positions in the collaboration network and how their influence might impact the overall dynamics. This network analysis provided a comprehensive view of the collaboration patterns among contributors, highlighting the roles of different demographic groups and the structure of the contributor community within the Rust ecosystem.
@@ 4. Results
Gender representation among contributors was consistently skewed across all projects, with individuals identified as women comprising only 3.4% in Rust, 3.7% in Rust Analyzer, and 3.8% in Cargo. These figures align with the survey’s report of 3.5% women participants (Team, 2024). However, gender disparity varied by country and project:
@@
Comparatively, the community distribution according to the 2023 Rust Annual Survey (Team, 2024), the top ten countries represented among participants were the United States (22%), Germany (12%), China (6%), the United Kingdom (6%), France (6%), Canada (3%), Russia (3%), the Netherlands (3%), Japan (3%), and Poland (3%), with participants hailing from approximately 100 countries.
Figure 2. Rust project contributors social network (A larger image is available in our replication package)

### v3

@@ Abstract.
Open source software communities thrive on global collaboration and contributions from diverse participants. This study explores the Rust programming language ecosystem to understand its contributors’ demographic composition and interaction patterns. Our objective is to investigate the phenomenon of participation inequality in key Rust projects and the presence of diversity among them. We studied GitHub pull request data from the year leading up to the release of the latest completed Rust community annual survey in 2023. Specifically, we extracted information from three leading repositories: Rust, Rust Analyzer, and Cargo, and used social network graphs to visualize the interactions and identify central contributors and subcommunities. Social network analysis has shown concerning disparities in gender and geographic representation among contributors who play pivotal roles in collaboration networks and the presence of varying diversity levels in the subcommunities formed. These results suggest that while the Rust community is globally active, the contributor base does not fully reflect the diversity of the wider user community. We conclude that there is a need for more inclusive practices to encourage broader participation and ensure that the contributor base aligns more closely with the diverse global community that utilizes Rust.
Repository Mining, Social networks, OSS communities, Rust, DEI
@@ 1. Introduction
Open source software (OSS) communities form the backbone of contemporary software development, enabling global collaboration and innovation. Participants from diverse backgrounds contribute in various ways (i.e., reporting issues, submitting pull requests (PRs), conducting code reviews, and updating documentation). These participants often contribute in their free time and at a level that aligns with their expertise and availability, fostering an environment of collective growth and learning (Lakhani and Wolf, 2005; Gerosa et al., 2021). Despite the seemingly vast pool of available developers, research shows that much of OSS development is driven by a small percentage of individuals. This phenomenon, known as “participation inequality” (Gasparini et al., 2020a), highlights the disparity in contribution levels, with most work concentrated among a few core contributors. Mature OSS communities, such as the Apache Software Foundation, [^1] https://www.apache.org/ are known to be inclusive (Feng et al., 2023) but have been plagued by the same issue (Chelkowski et al., 2016).
@@
Like many other open source software (OSS) projects, the Rust community is structured around a formal membership system (Toral et al., 2010). Sustained contributions to core Rust projects are crucial for the language’s evolution, and among these contributors, full members of the Rust GitHub organization occupy influential roles by virtue of their privileged access and responsibilities, such as voting on Requests for Comments (RFCs), which help shape the project’s strategic direction. [^3] https://forge.rust-lang.org/compiler/membership.html Despite their importance, the extent to which these contributors mirror the global diversity of Rust’s broader user base remains underexplored. This potential misalignment could significantly affect the community’s ability to foster an inclusive environment that attracts and retains a diverse range of contributors worldwide.
@@
  • RQ3: To what extent do subcommunities and top contributors within the Rust ecosystem exhibit diversity?
To answer these research questions, we focused on the core Rust compiler, Rust Analyzer (an integrated development environment (IDE) front-end), and Cargo (the Rust package manager). We examined the contributions and interactions within these repositories. Utilizing the GitHub API, we extracted PRs, authors, and reviewers. Then, we enriched this dataset with demographic information such as geographic location and likely gender. We then constructed social network graphs in R, applying the Louvain community detection algorithm to identify subcommunities within the network. Diversity within these subcommunities was quantitatively assessed using Blau’s Index. At the same time, various network metrics were calculated to evaluate the structure and dynamics of Rust contributor interactions.
Our findings revealed that the diversity of active contributors does not fully reflect the broader Rust user community’s demographic composition; this can be viewed in our anonymized replication that is publicly available (Dandamudi and Rodríguez-Pérez, 2024). The contributor base is heavily concentrated in specific geographic regions, with a predominance of contributors from the United States and Germany. Gender diversity remains extremely low, with men contributors dominating participation. Additionally, non-member contributors play a crucial role in the collaboration networks, often forming distinct clusters independent of official Rust organization members. Subcommunities demonstrate substantial geographic diversity; however, gender diversity remains notably low across all groups. These findings suggest that the community may need more inclusive practices to attract a broader range of participants to reflect its user base as it leads to more adoption, which in turn leads to more job creation that benefits the whole ecosystem (Carter and Groopman, 2021), aligns with Rust’s ethos of empowering everyone and capturing contributors from upcoming communities in Southern and Eastern parts of the world (GitHub, 2023).
@@ 2. Related work
The success and sustainability of OSS projects have been observed to correlate with their contributor communities’ collaborative efforts and network structures. Previous research has focused on understanding these dynamics, particularly through the lens of social network analysis and diversity studies.
@@ 2.1. Social Network Analysis in OSS Communities
An empirical analysis of GitHub project social team structures reveals a significant correlation between the speed of pull-request processing and the density of the collaboration network (Mezouar et al., 2019). The study found that smaller to medium-sized projects benefit from a tightly-knit core of maintainers who engage in repeated interactions, facilitating more efficient review and integration of contributions. Building on these findings, we aim to investigate similar dynamics within three significant repositories in the Rust ecosystem. Our analysis will examine the network structures between Rust Github organization members (typically those with privileged permissions and approval access) and non-members, further exploring how geographic and gender demographics influence interaction patterns.
@@ 2.3. Rust Community
Recent works have used network analysis (Imtiaz et al., 2023) in the Rust ecosystem, constructing developer networks based on collaboration activities. The study proposed integrating contributor reputation metrics to enhance dependency management practices, highlighting the importance of social factors in technical decision-making processes within OSS communities.
Our study seeks to bridge the current state of research done in the Rust ecosystem gap by explicitly analyzing the social dynamics and demographic diversity within the Rust community. By integrating social network analysis with demographic data such as geographic location and gender, we aim to uncover how diversity manifests within subcommunities and influences collaboration patterns. Unlike previous studies that have focused solely on centrality measures or diversity aspects, our research provides a comprehensive view combining both elements, offering deeper insights into the factors that drive participation and collaboration in the Rust ecosystem.
@@ 3.1. Projects Selection Criteria
As of January 2025, the Rust organization on GitHub comprises approximately 222 repositories. To focus our analysis on the most impactful projects, we selected the top three repositories based on the number of stars, a common indicator of popularity and community engagement (Ray et al., 2017). We excluded repositories, such as book, mdBook, and rustlings, as they are intended for documentation and learning purposes. After this refinement, our study concentrated on three core repositories:
@@
Furthermore, these three repositories represent distinct facets of the Rust ecosystem, each demanding unique skill sets and domain expertise. The Rust compiler involves low-level systems programming and compiler design, requiring contributors to understand language internals and performance optimization deeply. Rust Analyzer focuses on IDE integration and developer tooling, necessitating knowledge in language server protocols, code analysis, and user experience considerations. Cargo, as the package manager, revolves around dependency management and building systems, where contributors need proficiency in software distribution and project configuration. Each project follows unique development and collaboration practices tailored to the specific goals and challenges mentioned in their dedicated guidelines. [^5] https://rustc-dev-guide.rust-lang.org/contributing.html [^6] https://github.com/rust-lang/rust-analyzer/blob/master/docs/dev/guide.md [^7] https://doc.crates.io/contrib/ By including these repositories, our study captures a broad range of contribution dynamics and interactions within the Rust community, providing insights into how varying technical domains influence participation and collaboration dynamics. We classify contributors into two categories: members, accounts that belong to the public rust-lang GitHub organization and therefore have merge and branch-protection privileges, and non-members, who contribute without those elevated rights.
@@ 3.2. Extracting PRs, Author, and Reviewers
We posit that individuals who invested the time to complete the comprehensive Rust Annual Community Survey 2023 (The Rust Team, 2024), which takes approximately 25 minutes, represent the community’s distribution due to their significant level of engagement and commitment. Hence, to understand and compare the contributors’ dynamics preceding the Rust Annual Survey 2023, we collected PR data from the year prior to the survey’s release, specifically from December 18th, 2022, to December 18th, 2023.
@@
Additionally, we excluded PRs involving automated bots such as dependabot, renovatebot, and rustbot to ensure our analysis focused solely on human contributors.
3.3. Enriching Data with Full Name, Location, and Membership
@@
To ascertain the geographical distribution of contributors, we analyzed the location details provided in user profiles. We utilized the geograpy3888https://pypi.org/project/geograpy3/ Python package, which leverages the Natural Language Toolkit (NLTK) framework, to parse the city or location strings and extract corresponding country information. The extracted data were then converted to ISO2 country codes [^9] https://www.iso.org/obp/ui/#search using the country_converter package. To ensure the highest accuracy, we manually verified the location data by searching the location details (city, country emoji, etc) online with the assigned country code, correcting any inconsistencies or errors identified during the automated extraction process. Overall, about a third of the data was inconsistently classified. For example, “Earth” and “UK” were classified under the “US” country code.
Table 1. Statistics of code contributors’ demographics in three crucial Rust projects
@@
  • Values shown in bold show higher likely women participation, full non-member participation, and low contributor share.
For gender identification, we relied on contributors’ full names. We employed Stanford’s Named Entity Recognition (NER) tool, available through the Python package stanza [^10] https://stanfordnlp.github.io/stanza/ to determine whether the extracted names referred to persons. This filtered out usernames that might represent organizations, projects, or non-person entities. Only names recognized as “PERSON” entities, with additional manual verification, were considered in subsequent analyses. Subsequently, we used the NamSor API, a widely recognized name classification tool in academic research (Sebo, 2021; Qiu et al., 2019), to predict the likely gender of each contributor. By inputting the full name and corresponding ISO2 country code into NamSor, we obtained gender probabilities for each individual. To maintain a high level of confidence in our gender assignments, we included only those contributors for whom NamSor provided a gender probability of 80% or higher. Any cases with lower confidence levels or potential false positives underwent additional manual verification. Both authors independently reviewed these ambiguous cases, and only upon mutual agreement were the final gender classifications assigned.
@@ 3.4. Social Network Analysis
Post enriched data collection, we constructed social network graphs using the R programming language and the igraph [^11] https://igraph.org/ package. Each contributor is represented as a node, and the edges represent user interactions in these graphs. We developed two types of graphs: directed and undirected. In the directed graphs, where edges from authors to reviewers represent reviewed pull requests (PRs), and undirected graphs, where edges denote PR-based interactions. The former captured reciprocity and influence, while the latter revealed community structures, offering a holistic view of contributor dynamics.
To identify clusters within the social network, we applied the Louvain community detection algorithm, which efficiently uncovers communities by optimizing modularity (Blondel et al., 2008). This algorithm allowed us to detect subcommunities where contributors are more densely connected among themselves than with the rest of the network. Incorporating demographic details into our analysis enabled us to understand various collaboration dynamics. We examined how members and non-members of the Rust GitHub organization collaborate, exploring whether official affiliation influences interaction patterns. Additionally, we assessed the diversity within sub-communities by considering contributors’ geographic locations (using ISO2 country codes) and likely gender. This approach provided insights into how contributors from different backgrounds interact within the network.
The graphs were constructed by including nodes only when the respective node demographic information was available. To ensure the accuracy and relevance of our network analysis, we removed self-loops, which are edges that connect nodes to themselves, as they do not represent meaningful collaborations between different contributors. Furthermore, we excluded isolated nodes that remained disconnected after the removal of self-loops, focusing our analysis on the interconnected portion of the network where actual interactions occur.
@@
To capture subcommunity geographic diversity, we constructed graphs using ISO2 country codes and used Blau’s Index(Blau, 1977) to calculate the diversity quantitatively in identified subcommunities across the projects (Shameer et al., 2023).
@@
A higher Blau’s Index value indicates greater diversity within the subcommunity, while a lower value suggests homogeneity. By applying this metric to the geographic distributions of contributors, we could quantitatively compare the diversity levels across different subcommunities within the Rust ecosystem. Additionally, we employed centrality measures (Newman, 2018), such as degree centrality (the number of direct connections a node has), betweenness centrality (the extent to which a node lies on paths between other nodes), and eigenvector centrality (the influence of a node based on the importance of its connections), to identify influential contributors within the network. By examining the demographics of these central contributors, we aimed to understand which user groups hold significant positions in the collaboration network and how their influence might impact the overall dynamics. This network analysis provided a comprehensive view of the collaboration patterns among contributors, highlighting the roles of different demographic groups and the structure of the contributor community within the Rust ecosystem.
@@ 4. Results
Gender representation among contributors was consistently skewed across all projects, with individuals identified as women comprising only 3.4% in Rust, 3.7% in Rust Analyzer, and 3.8% in Cargo. These figures align with the survey’s report of 3.5% women participants (The Rust Team, 2024). However, gender disparity varied by country and project:
@@
Comparatively, the community distribution according to the 2023 Rust Annual Survey (The Rust Team, 2024), the top ten countries represented among participants were the United States (22%), Germany (12%), China (6%), the United Kingdom (6%), France (6%), Canada (3%), Russia (3%), the Netherlands (3%), Japan (3%), and Poland (3%), with participants hailing from approximately 100 countries.
Figure 2. Rust project contributors’ social network (A larger image is available in our replication package)